OpenStudy (anonymous):

Identify cos^2 theta(1+tan^2 theta)=1. Start with the left-hand side.(Hint: There should be 3 step.)

Parth (parthkohli):

Do you know any Pythagorean identity associated with \(1 + \tan^2 \theta\)?

OpenStudy (anonymous):

\[\sec^2 \theta\] right?

Parth (parthkohli):

That's right! What is \(\cos^2 \theta \times \sec^2 \theta\)?

OpenStudy (anonymous):

0?

Parth (parthkohli):

\[\cos^2\theta \times \dfrac{1}{\cos^2\theta}\]

OpenStudy (anonymous):

I figured it out, but thank you for your help! :)